# A Gentle Introduction to the Axiom of Choice

### v2

@@ 3.2 Cracking walnuts with a sledgehammer
For the next result, we introduce the following notation: given a set $X$, let $\mathcal{W}(X)$ be the set of all well-orderable subsets of $X$, i.e. those subsets of $X$ that can be well-ordered (see footnote 1).
@@
Recall that $\mathsf{AC}$ is equivalent to the well-ordering principle. So, over $\mathsf{ZF}$—where some sets may admit no well-ordering—this is a strengthening of Cantor’s theorem, which states that $\lvert X\rvert<\lvert\mathcal{P}(X)\rvert$, where $\mathcal{P}(X)$ denotes the power set of $X$, i.e. the set of all subsets of $X$. By contrast, assuming $\mathsf{AC}$, $\mathcal{W}(X)=\mathcal{P}(X)$, so this result follows from Cantor’s theorem.
@@
Cantor-Schröder-Bernstein Theorem ($\mathsf{CSB}$). For any sets $X$ and $Y$, if $\lvert X\rvert\leqslant\lvert Y\rvert$ and $\lvert Y\rvert\leqslant\lvert X\rvert$, then $\lvert X\rvert=\lvert Y\rvert$. [^18] This result also has its own extremely interesting history, which is discussed in detail in [Hin13]. In particular, a more accurate title for the theorem would also include the names of Dedekind and Zermelo.
Over $\mathsf{ZF}$, the proof is fairly hands-on. Given injections $f:X\to Y$ and $g:Y\to X$, rather than merely proving the existence of a bijection, one must construct a bijection $h:X\to Y$ by stringing together these maps (see, for example, [Jec97]). Assuming $\mathsf{AC}$, the sets $X$ and $Y$ are well-orderable. In this case, standard results about well-orderings provide a “conceptually simpler” proof. [^19] In this case, the “simplicity” of the $\mathsf{AC}$-proof of $\mathsf{CSB}$ depends on taking for granted standard facts about well-orderings, so it may not necessarily be mathematically simpler. For instance, this argument relies on the trichotomy of well-orderings, i.e. given two well-orderings $W_{1}$ and $W_{2}$, either $W_{1}\cong W_{2}$ or one is isomorphic to an initial segment of the other, and on the fact that if $W$ is a well-ordered set and $V$ is a subset of $W$ with the induced ordering, then $W$ is not isomorphic to a proper initial segment of $V$. In contrast, the $\mathsf{ZF}$ proof, though tedious, does not require this additional machinery. A discussion by Hamkins and Schweber on whether or not this proof is actually simpler is in the comments of Hamkins’ answer in [Ham23].
@@
A standard alternative is the Galvin-Glazer argument, described by Hindman as “the shortest and, to my mind, prettiest proof” of Hindman’s theorem [Hin79]. Formalizing this proof requires several more levels of the cumulative hierarchy, and this proof also relies on two applications of $\mathsf{AC}$. However, in contrast to the original, this argument is one that can reasonably be remembered. Since this proof is still outside of the scope of this article, we provide a brief outline of the applications of $\mathsf{AC}$ in this proof, while omitting details. The detailed proof can be found in [Hin79].
@@ 3.3 Because you can!
The answers to both questions are on the side of consistency. That is, each of $\mathsf{ZFC}$ and $\mathsf{ZF}+\neg\mathsf{AC}$ is consistent. But to prove these answers we need an additional assumption, namely that $\mathsf{ZF}$ (without any commitment for or against $\mathsf{AC}$) is consistent. The additional assumption is required because of Gödel’s second incompleteness theorem [Göd31], which says that a reasonable202020“Reasonable” means that (1) the theory is consistent, (2) it can prove basic facts about the arithmetic of natural numbers, and (3) its axioms can be listed by an algorithm. In particular, $\mathsf{ZFC}$ is reasonable. theory cannot prove its own consistency. Thus, what is actually proved is that if $\mathsf{ZF}$ is consistent then so are $\mathsf{ZFC}$ and $\mathsf{ZF}+\neg\mathsf{AC}$. In other words, if $\mathsf{ZF}$ were to become inconsistent when one adds $\mathsf{AC}$ (or $\neg\mathsf{AC}$) as a new axiom, then $\mathsf{ZF}$ itself was already inconsistent; the inconsistency is not the fault of the added axiom.
@@
This description of the cumulative hierarchy involves vague notions, “all sets” and “transfinitely forever,” so it cannot be used directly as a basis for mathematical proofs. Despite the vagueness, the intuition of the cumulative hierarchy supports some precise statements about sets, and a useful collection of those was isolated in an axiomatic system $\mathsf{ZFA}$ (meaning $\mathsf{ZF}$ with atoms, and $\mathsf{ZFCA}$ adds $\mathsf{AC}$). $\mathsf{ZF}$ is the special case where there are no atoms. [^21] Even without atoms, there are plenty of sets: $\varnothing$ at level 1, $\{\varnothing\}$ at level 2, $\{\{\varnothing\}\}$ and $\{\varnothing,\{\varnothing\}\}$ at level 3, and so on. Infinite sets appear at the transfinite levels. It turns out that atoms are never really needed in mathematics, so $\mathsf{ZFC}$, rather than $\mathsf{ZFCA}$, has become the widely accepted axiomatic basis for set theory and thus for the rest of mathematics.
@@
The first phase began in 1922 with Abraham Fraenkel’s proof that $\mathsf{AC}$ cannot be proved in $\mathsf{ZFA}$ [Fra22]. He began with a universe satisfying $\mathsf{ZFA}$ and having infinitely many atoms, and he defined a sub-universe $S$ of “sufficiently symmetric” [^23] Here, “symmetry” refers to behavior when the atoms are arbitrarily permuted. Such a permutation $\pi$ produces a permutation of $V$. (In fact, $\pi$ produces an automorphism of $V$.) To apply $\pi$ to a set, simply apply $\pi$ to all its elements. Since the elements occur earlier in the cumulative hierarchy, this is a legitimate inductive definition. A set or atom $x$ is “sufficiently symmetric” if there is a finite set $E\subseteq A$ such that all permutations that fix all elements of $E$ also fix $x$. (It is tempting to simplify this definition by requiring $x$ to be fixed by all permutations. Unfortunately these “completely symmetric” sets don’t satisfy $\mathsf{ZFA}$. For example, both $\varnothing$ and $A$ are completely symmetric but they are unequal despite having the same completely symmetric members, namely none.) sets to violate $\mathsf{AC}$—for example, the set of atoms admits no symmetric well-ordering (or even linear ordering)—while satisfying all of the $\mathsf{ZFA}$ axioms. Fraenkel’s sub-universe consists of the hereditarily sufficiently symmetric sets and atoms, i.e., those $x$ such that $x$ and all its members, and all their members, $\dots$ are sufficiently symmetric. It is fairly easy to see that $\mathsf{AC}$ is false in this sub-universe. For example, all the atoms and the set $A$ of atoms are sufficiently symmetric, but no linear ordering of $A$ (let alone a well-ordering) is sufficiently symmetric. More work is needed to show that this sub-universe satisfies all the $\mathsf{ZFA}$ axioms [Jec73].
Fraenkel’s method was extended, especially by Andrzej Mostowski, to obtain considerable information about non-implications between various consequences of $\mathsf{AC}$ [Fra37, Mos39]. For example, even if one adds to $\mathsf{ZFA}$ an axiom saying that every set can be linearly ordered, one still cannot deduce the axiom of choice. Later, Ernst Specker systematized the theory of such permutation models (or Fraenkel-Mostowski models or Fraenkel-Mostowski-Specker models) in terms of an arbitrary group of permutations of atoms (to define “symmetric”) and an arbitrary normal filter of subgroups (to define “sufficiently”) [Spe57]. For more information about permutation models of $\mathsf{ZFA}$, see [Jec73, Chapter 4].

### v3

@@ 3.2 Cracking walnuts with a sledgehammer
For the next result, we introduce the following notation: given a set $X$, let $\mathcal{W}(X)$ be the set of all well-orderable subsets of $X$, i.e., those subsets of $X$ that can be well-ordered (see footnote 1).
@@
Recall that $\mathsf{AC}$ is equivalent to the well-ordering principle. So, over $\mathsf{ZF}$—where some sets may admit no well-ordering—this is a strengthening of Cantor’s theorem, which states that $\lvert X\rvert<\lvert\mathcal{P}(X)\rvert$, where $\mathcal{P}(X)$ denotes the power set of $X$, i.e., the set of all subsets of $X$. By contrast, assuming $\mathsf{AC}$, $\mathcal{W}(X)=\mathcal{P}(X)$, so this result follows from Cantor’s theorem.
@@
Cantor–Schröder–Bernstein Theorem ($\mathsf{CSB}$). For any sets $X$ and $Y$, if $\lvert X\rvert\leqslant\lvert Y\rvert$ and $\lvert Y\rvert\leqslant\lvert X\rvert$, then $\lvert X\rvert=\lvert Y\rvert$. [^21] This result also has its own extremely interesting history, which is discussed in detail in [Hin13]. In particular, a more accurate title for the theorem would also include the names of Dedekind and Zermelo.
Over $\mathsf{ZF}$, the proof is fairly hands-on. Given injections $f:X\to Y$ and $g:Y\to X$, rather than merely proving the existence of a bijection, one must construct a bijection $h:X\to Y$ by stringing together these maps (see, for example, [Jec97]). Assuming $\mathsf{AC}$, the sets $X$ and $Y$ are well-orderable. In this case, standard results about well-orderings provide a “conceptually simpler” proof. [^22] In this case, the “simplicity” of the $\mathsf{AC}$-proof of $\mathsf{CSB}$ depends on taking for granted standard facts about well-orderings, so it may not necessarily be mathematically simpler. For instance, this argument relies on the trichotomy of well-orderings, i.e., given two well-orderings $W_{1}$ and $W_{2}$, either $W_{1}\cong W_{2}$ or one is isomorphic to an initial segment of the other, and on the fact that if $W$ is a well-ordered set and $V$ is a subset of $W$ with the induced ordering, then $W$ is not isomorphic to a proper initial segment of $V$. In contrast, the $\mathsf{ZF}$ proof, though tedious, does not require this additional machinery. A discussion by Hamkins and Schweber on whether or not this proof is actually simpler is in the comments of Hamkins’ answer in [Ham23].
@@
A standard alternative is the Galvin–Glazer argument, described by Hindman as “the shortest and, to my mind, prettiest proof ” of Hindman’s theorem [Hin79]. Formalizing this proof requires several more levels of the cumulative hierarchy, and this proof also relies on two applications of $\mathsf{AC}$. However, in contrast to the original, this argument is one that can reasonably be remembered. Since this proof is still outside of the scope of this article, we provide a brief outline of the applications of $\mathsf{AC}$ in this proof, while omitting details. The detailed proof can be found in [Hin79].
@@ 3.3 Because you can!
The answers to both questions are on the side of consistency. That is, each of $\mathsf{ZFC}$ and $\mathsf{ZF}+\neg\mathsf{AC}$ is consistent. But to prove these answers we need an additional assumption, namely that $\mathsf{ZF}$ (without any commitment for or against $\mathsf{AC}$) is consistent. The additional assumption is required because of Gödel’s second incompleteness theorem [Göd31], which says that a reasonable232323“Reasonable” means that (1) the theory is consistent, (2) it can prove basic facts about the arithmetic of natural numbers, and (3) its axioms can be listed by an algorithm. In particular, $\mathsf{ZFC}$ is reasonable. theory cannot prove its own consistency. Thus, what is actually proved is that if $\mathsf{ZF}$ is consistent then so are $\mathsf{ZFC}$ and $\mathsf{ZF}+\neg\mathsf{AC}$. In other words, if $\mathsf{ZF}$ were to become inconsistent when one adds $\mathsf{AC}$ (or $\neg\mathsf{AC}$) as a new axiom, then $\mathsf{ZF}$ itself was already inconsistent; the inconsistency is not the fault of the added axiom.
@@
This description of the cumulative hierarchy involves vague notions, “all sets” and “transfinitely forever,” so it cannot be used directly as a basis for mathematical proofs. Despite the vagueness, the intuition of the cumulative hierarchy supports some precise statements about sets, and a useful collection of those was isolated in an axiomatic system $\mathsf{ZFA}$ (meaning $\mathsf{ZF}$ with atoms, and $\mathsf{ZFCA}$ adds $\mathsf{AC}$). $\mathsf{ZF}$ is equivalent to $\mathsf{ZFA}$ plus the axiom that there are no atoms [^24] Even without atoms, there are plenty of sets: $\varnothing$ at level 1, $\{\varnothing\}$ at level 2, $\{\{\varnothing\}\}$ and $\{\varnothing,\{\varnothing\}\}$ at level 3, and so on. Infinite sets appear at the transfinite levels., so $\mathsf{ZFA}$ is consistent if $\mathsf{ZF}$ is. More importantly for our purposes, the consistency of $\mathsf{ZF}$ implies the consistency of $\mathsf{ZFA}$ plus the axiom that there are infinitely many atoms [Jec73, Problem 4.6.1.]. It turns out that atoms are never really needed in mathematics, so $\mathsf{ZFC}$, rather than $\mathsf{ZFCA}$, has become the widely accepted axiomatic basis for set theory and thus for the rest of mathematics.
@@
The first phase began in 1922 with Abraham Fraenkel’s proof that $\mathsf{AC}$ cannot be proved in $\mathsf{ZFA}$ [Fra22]. He began with a universe satisfying $\mathsf{ZFA}$ and having infinitely many atoms, and he defined a sub-universe $S$ of “sufficiently symmetric” [^26] Here, “symmetry” refers to behavior when the atoms are arbitrarily permuted. Such a permutation $\pi$ produces a permutation of $V$. (In fact, $\pi$ produces an automorphism of $V$.) To apply $\pi$ to a set, simply apply $\pi$ to all its elements. Since the elements occur earlier in the cumulative hierarchy, this is a legitimate inductive definition. A set or atom $x$ is “sufficiently symmetric” if there is a finite set $E\subseteq A$ such that all permutations that fix all elements of $E$ also fix $x$. (It is tempting to simplify this definition by requiring $x$ to be fixed by all permutations. Unfortunately these “completely symmetric” sets don’t satisfy $\mathsf{ZFA}$. For example, both $\varnothing$ and $A$ are completely symmetric but they are unequal despite having the same completely symmetric members, namely none.) sets to violate $\mathsf{AC}$—for example, the set of atoms admits no symmetric well-ordering (or even linear ordering)—while satisfying all of the $\mathsf{ZFA}$ axioms. Fraenkel’s sub-universe consists of the hereditarily sufficiently symmetric sets and atoms, i.e., those $x$ such that $x$ and all its members, and all their members, $\dots$ are sufficiently symmetric. It is fairly easy to see that $\mathsf{AC}$ is false in this sub-universe. For example, all the atoms and the set $A$ of atoms are sufficiently symmetric, but no linear ordering of $A$ (let alone a well-ordering) is sufficiently symmetric. More work is needed to show that this sub-universe satisfies all the $\mathsf{ZFA}$ axioms; see [Jec73] or [Hal25].
Fraenkel’s method was extended, especially by Andrzej Mostowski, to obtain considerable information about non-implications between various consequences of $\mathsf{AC}$ [Fra37, Mos39]. For example, even if one adds to $\mathsf{ZFA}$ an axiom saying that every set can be linearly ordered, one still cannot deduce the axiom of choice. Later, Ernst Specker systematized the theory of such permutation models (or Fraenkel–Mostowski models or Fraenkel–Mostowski–Specker models) in terms of an arbitrary group of permutations of atoms (to define “symmetric”) and an arbitrary normal filter of subgroups (to define “sufficiently”) [Spe57]. For more information about permutation models of $\mathsf{ZFA}$, see [Jec73, Chapter 4].